# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

### v1

@@ Abstract
Video Frame Interpolation (VFI) is a fundamental Low-Level Vision (LLV) task that synthesizes intermediate frames between existing ones while maintaining spatial and temporal coherence. VFI techniques have evolved from classical motion compensation-based approach to deep learning-based approach, including kernel-, flow-, hybrid-, phase-, GAN-, Transformer-, Mamba-, and more recently diffusion model-based approach. We introduce AceVFI, the most comprehensive survey on VFI to date, covering over 250+ papers across these approaches. We systematically organize and describe VFI methodologies, detailing the core principles, design assumptions, and technical characteristics of each approach. We categorize the learning paradigm of VFI methods namely, Center-Time Frame Interpolation (CTFI) and Arbitrary-Time Frame Interpolation (ATFI). We analyze key challenges of VFI such as large motion, occlusion, lighting variation, and non-linear motion. In addition, we review standard datasets, loss functions, evaluation metrics. We examine applications of VFI including event-based, cartoon, medical image VFI and joint VFI with other LLV tasks. We conclude by outlining promising future research directions to support continued progress in the field. This survey aims to serve as a unified reference for both newcomers and experts seeking a deep understanding of modern VFI landscapes. We maintain an up-to-date project page: https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation.
@@
Video Frame Interpolation, Generative Inbetweening, Video Generation, Low-Level Vision
publicationid: pubid:
1 Introduction
Video Frame Interpolation (VFI) aims to increase the temporal resolution (i.e., frame rate) of a video sequence by synthesizing one or more intermediate frames between given consecutive frames. This task serves a broad range of applications, including novel view synthesis [1, 2, 3, 4], slow-motion generation [5, 6, 7, 8, 9, 10], video compression [11, 12, 13, 14], video prediction [15, 16, 13, 17], and diverse generation tasks such as co-speech reenactment [18], human motion synthesis [19], and facial animation [20]. A key advantage of VFI lies in its ability to synthesize perceptually smooth and temporally coherent motion, aligning well with the temporal characteristics of the human visual system (HVS). High-frame-rate (HFR) content reduces artifacts such as motion blur and judder [21, 22], thereby enhancing the visual quality in high-resolution (HR) and immersive media. This makes VFI particularly valuable in latency-sensitive and fidelity-critical scenarios such as sports broadcasting, interactive gaming, and virtual reality. Finally, in streaming pipelines, VFI also enables bandwidth-efficient video transmission by reconstructing intermediate frames locally, reducing the need to transmit full frame sequences [21].
Figure 1: General process of VFI. Given two input frames $I_{0}$ and $I_{1}$, the VFI model $\mathcal{F}$ synthesizes one or more intermediate frames. ×$\times n$ interpolation denotes synthesizing $n{-}1$ intermediate frames to increase the frame rate by a factor of $n$.
Formally, given two frames $I_{0}$ and $I_{1}$, a VFI model $\mathcal{F}$ estimates the interpolated frame $\hat{I}_{t}$ at time $t\in(0,1)$:
@@ 1 Introduction
As shown in Fig. 1, interpolating $n{-}1$ frames between each input pair increases the frame rate by a factor of $n$. For example, generating seven intermediate frames per interval transforms a 30fps video into 240fps.
1.1 General Pipeline of VFI
@@
Figure 2: General pipeline of VFI. Given two input frames $I_{0}$ and $I_{1}$, the feature representations $F_{0}$ and $F_{1}$ are extracted, which are then aligned to the target time $t$ using the estimated motion. The temporal alignment can be performed on either input RGB pixels or features, resulting in $\hat{I}_{0\rightarrow t},\hat{I}_{1\rightarrow t}$ or $\hat{F}_{0\rightarrow t},\hat{F}_{1\rightarrow t}$. Finally, a Frame Synthesis module blends the aligned inputs to produce the interpolated frame $\hat{I}_{t}$. This pipeline highlights the four core stages of VFI: Feature Extraction, Motion Estimation, Temporal Alignment, and Frame Synthesis.
@@
As shown in Fig. 2, the general VFI pipeline consists of four stages. (i) Feature Extraction: Input frames $I_{0}$ and $I_{1}$ are first passed through a feature extraction network [23, 24, 25] to obtain deep features $F_{0}$ and $F_{1}$. These features encode spatial and semantic information suitable for subsequent motion reasoning [26]. (ii) Motion Estimation: The temporal correspondence, commonly referred to as motion, is estimated. Motion estimation is performed either explicitly (e.g., optical flow [27]) or implicitly via kernels [28, 29, 16, 30, 31, 32, 33, 34, 35, 36, 37, 8, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], phase [48, 49], attention maps [50, 51, 52, 53, 54, 55, 56, 57], or cost volumes [35, 58, 59, 53]. (iii) Temporal Alignment: The estimated motion is used to temporally align the input pixels or features to the target time $t$, resulting in $\hat{I}_{0\rightarrow t}$, $\hat{I}_{1\rightarrow t}$ or $\hat{F}_{0\rightarrow t}$, $\hat{F}_{1\rightarrow t}$. There are four types of alignment strategies. Kernel-based alignment (Fig. 3 (a)) aggregates local or non-local information from the inputs using learned, spatially-adaptive kernels. These kernels implicitly encode motion by adapting their spatial weights based on local context, allowing motion-aware alignment without explicit flow estimation. Flow-based alignment (Fig. 3 (b)) warps inputs guided by the estimated flow. Forward warping [60] maps source pixels (i.e., input pixels) to their estimated locations in the target frame. Backward warping [61] samples from the source based on coordinates in the target frame, effectively pulling information from the source toward the desired time. Attention-based alignment (Fig. 3 (c)) replaces explicit geometric warping with attention-weighted aggregation [51, 53]. By computing soft correspondences between elements across input frames, this approach can adaptively focus on semantically relevant regions and align contents even across large spatial-temporal gaps. Cost volume-based alignment (Fig. 3 (d)) constructs dense similarity volumes between feature maps, enabling precise correspondence modeling across space and time. (iv) Frame Synthesis: Finally, the aligned inputs are blended to synthesize the final interpolated frame using either simple averaging, weighted blending or synthesis networks [62].
Figure 3: Temporal alignment process. The input frames $(I_{0},I_{1})$ or their features $(F_{0},F_{1})$ are temporally aligned toward a target time $t$. (a) kernel-based alignment using spatially-adaptive filters, (b) flow-based alignment guided by optical flow, (c) attention-based alignment using attention-weighted correspondences, and (d) cost volume-based alignment through pixel- or feature-level similarity estimation.
{forest}
[…]
Combined (Sec. 2.2.3), color=colorSec2!95, fill=colorSec2!25, text=black] [Phase-based (Sec. 2.2.4), color=colorSec2, fill=colorSec2!25, text=black] [GAN-based (Sec. 2.2.5), color=colorSec2, fill=colorSec2!25, text=black] [Transformer-based (Sec. 2.2.6), color=colorSec2, fill=colorSec2!25, text=black] [Mamba-based (Sec. 2.2.7), color=colorSec2, fill=colorSec2!25, text=black] ] [Diffusion-based (Sec. 2.3), color=colorSec2, fill=colorSec2!50, text=black] ] [III. Learning
Paradigm (Sec. 3), color=colorSec3, fill=colorSec3, text=black [Center-Time Frame Interpolation
(CTFI) (Sec. 3.1), color=colorSec3, fill=colorSec3!50, text=black] [Arbitrary-Time Frame Interpolation
(ATFI) (Sec. 3.2), color=colorSec3, fill=colorSec3!50, text=black] [Training Strategy (Sec. 3.3), color=colorSec3!95, fill=colorSec3!30, text=black [CTFI Training Strategy
(CTFI-TS) (Sec. 3.3.1), color=colorSec3, fill=colorSec3!25, text=black] [ATFI Training Strategy
(ATFI-TS) (Sec. 3.3.2), color=colorSec3, fill=colorSec3!25, text=black] ] [Loss Functions (Sec. 3.4), color=colorSec3, fill=colorSec3!50, text=black [Reconstruction Loss (Sec. 3.4.1), color=colorSec3, fill=colorSec3!25, text=black] [Perceptual Loss (Sec. 3.4.2), color=colorSec3, fill=colorSec3!25, text=black] [Adversarial Loss (Sec. 3.4.3), color=colorSec3, fill=colorSec3!25, text=black] [Flow Loss (Sec. 3.4.4), color=colorSec3, fill=colorSec3!25, text=black] ] ] [IV. VFI Challenges (Sec. 4), color=colorSec4, fill=colorSec4, text=black [Large Motion (Sec. 4.1), color=colorSec4, fill=colorSec4!50, text=black] [Occlusion (Sec. 4.2), color=colorSec4, fill=colorSec4!50, text=black] [Lighting Variation (Sec. 4.3), color=colorSec4, fill=colorSec4!50, text=black] [Non-linear Motion (Sec. 4.4), color=colorSec4, fill=colorSec4!50, text=black] ] [V. Datasets and
Evaluation Metrics (Sec. 5), color=colorSec5, fill=colorSec5, text=black [Datasets (Sec. 5.1), color=colorSec5, fill=colorSec5!50, text=black [Triplet Datasets (Sec. 5.1.1), color=colorSec5, fill=colorSec5!25, text=black] [Multi-frame Datasets (Sec. 5.1.2), color=colorSec5, fill=colorSec5!25, text=black] ] [Data Augmentation (Sec. 5.2), color=colorSec5, fill=colorSec5!50, text=black] [Evaluation Metrics (Sec. 5.3), color=colorSec5, fill=colorSec5!50, text=black [Image-level Metrics (Sec. 5.3.1), color=colorSec5, fill=colorSec5!25, text=black] [Perceptual Metrics (Sec. 5.3.2), color=colorSec5, fill=colorSec5!25, text=black] [Video-level Metrics (Sec. 5.3.3), color=colorSec5, fill=colorSec5!25, text=black] ] ] [VI. Applications (Sec. 6), color=colorSec6, fill=colorSec6, text=black [Event-based VFI (Sec. 6.1), color=colorSec6, fill=colorSec6!50, text=black] [Cartoon VFI (Sec. 6.2), color=colorSec6, fill=colorSec6!50, text=black] [Medical Image VFI (Sec. 6.3), color=colorSec6, fill=colorSec6!50, text=black] [Joint Task VFI (Sec. 6.4), color=colorSec6, fill=colorSec6!50, text=black] ] [VII. Future Research
Directions (Sec. 7), color=colorSec7, fill=colorSec7, text=black [Video Streaming Service (Sec. 7.1), color=colorSec7, fill=colorSec7!50, text=black] [All-in-One
LLV Video Restoration (Sec. 7.2), color=colorSec7, fill=colorSec7!50, text=black] [3D and 4D
Scene Understanding (Sec. 7.3), color=colorSec7, fill=colorSec7!50, text=black] ] ]
Figure 4: Overview of the survey structure. The figure summarizes the hierarchical organization of the survey, including methodological categories (Sec. 2), learning paradigms (Sec. 3), key challenges (Sec. 4), datasets and evaluation metrics (Sec. 5), applications (Sec. 6), and future research directions (Sec. 7).
1.2 Methodology Overview
VFI methodologies can be broadly classified into three major categories: motion compensation-based [63, 64, 65, 66, 67, 68, 69, 70, 71, 72], deep learning-based [61, 16, 73, 7, 74, 75, 76, 62, 60, 77, 78, 79, 80, 81, 58, 59, 82, 83, 84, 55, 85, 54, 53, 86, 87, 88, 89, 90, 91, 92, 28, 29, 32, 31, 50, 34, 39, 40, 41, 42, 43, 44, 45, 46, 30, 33, 35, 36, 37, 38, 93, 94, 95, 96, 74, 97, 48, 49, 51, 52, 56, 57, 98, 99, 100], and diffusion models (DMs)-based approach [101, 96, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119].
The motion compensation-based approach dominated the pre-deep-learning era, offering a straightforward two-stage strategy: estimating motion explicitly and warping frames accordingly. While effective under simple motion, its reliance on hand-crafted rules and block-based assumptions limits its ability to handle occlusions and complex, non-rigid dynamics.
With the advent of convolutional neural networks (CNNs) [120], the field shifted toward deep learning-based approach. This approach replaces heuristic pipelines with end-to-end frameworks that learn motion patterns and appearance features directly from data. As a result, they significantly improve robustness under diverse and challenging conditions. Further methodological details are discussed in Sec. 2.2.
More recently, DMs have been introduced as a generative perspective to VFI, framing it as a conditional denoising process rather than a deterministic prediction task. This has expanded the scope of VFI into the broader paradigm of Generative Inbetweening [112, 111], enabling uncertainty-aware interpolation and semantically diverse frame synthesis. This shift not only enhances robustness in ambiguous motion scenarios but also opens the door to multi-modal guidance (e.g., text, depth, or motion priors), redefining the role of VFI in creative and interactive video generation.
Overview. Fig. 4 shows the overall structure of this paper. Sec. 2 analyzes methodological taxonomies of VFI. Sec. 3 introduces and compares the two principal learning paradigms of VFI, and further examines their corresponding training strategies and loss functions. Sec. 4 discusses major challenges in VFI, along with how recent methods address them. Sec. 5 reviews common datasets and evaluation metrics. Sec. 6 explores applications of VFI across diverse domains. Finally, Sec. 7 presents future research directions of VFI.
2 Methodology
2.1 Motion Compensation-based

### v3

# AceVFI: A Comprehensive Survey of Advances in Video Frame Interpolation

@@ Abstract
Video Frame Interpolation (VFI) is a core low-level vision task that synthesizes intermediate frames between existing ones while ensuring spatial and temporal coherence. Over the past decades, VFI methodologies have evolved from classical motion compensation-based approach to a wide spectrum of deep learning-based approaches, including kernel-, flow-, hybrid-, phase-, GAN-, Transformer-, Mamba-, and most recently, diffusion-based models. We introduce AceVFI, a comprehensive and up-to-date review of the VFI field, covering over 250 representative papers. We systematically categorize VFI methods based on their core design principles and architectural characteristics. Further, we classify them into two major learning paradigms: Center-Time Frame Interpolation (CTFI) and Arbitrary-Time Frame Interpolation (ATFI). We analyze key challenges in VFI, including large motion, occlusion, lighting variation, and non-linear motion. In addition, we review standard datasets, loss functions, evaluation metrics. We also explore VFI applications in other domains and highlight future research directions. This survey aims to serve as a valuable reference for researchers and practitioners seeking a thorough understanding of the modern VFI landscape. We maintain an up-to-date project page: https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation.
@@
Video Frame Interpolation, Generative Inbetweening, Video Generation, Low-Level Vision.
[Uncaptioned image][Uncaptioned image] [Uncaptioned image] AceVFI: A Comprehensive Survey of Advances
in Video Frame Interpolation
Dahyeon Kye1    Changhyun Roh1    Sukhun Ko1    Chanho Eom2    Jihyong Oh1,† 1CMLab, Chung-Ang University    2Perceptual AI LAB, Chung-Ang University {rpekgus, changhyunroh, looloo330, cheom, jihyongoh}@cau.ac.kr https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation
footnotetext: Copyright © 2026 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.22footnotetext: Corresponding author.
I Introduction
Video Frame Interpolation (VFI) aims to increase the temporal resolution (i.e., frame rate) of a video sequence by synthesizing one or more intermediate frames between given consecutive frames. This task serves a broad range of applications, including novel view synthesis [214, 284, 60, 134], slow-motion generation [107, 252, 6, 244, 103, 92], video compression [238, 38, 106, 215], video prediction [154, 148, 106, 81], and diverse generation tasks such as co-speech reenactment [141], human motion synthesis [140], and facial animation [13]. In many of these scenarios, VFI is not merely an optional post-processing tool but a practically irreplaceable component: for slow-motion generation, high-frame-rate (HFR) capture typically requires specialized sensors, strong illumination, and large bandwidth or storage, which are often unavailable in consumer or legacy footage, so once a scene has been recorded at a low-frame-rate (LFR), additional real frames cannot be acquired retrospectively, so learned VFI becomes the only viable way to synthesize plausible in-between frames [153]. Likewise, in novel view synthesis and 4D scene reconstruction, densely sampling both viewpoints and time is often infeasible due to camera cost, calibration and synchronization overhead, and storage constraints, so VFI is used to temporally densify sparse input sequences, providing smoother motion trajectories and more continuous temporal coverage for downstream 3D/4D methods [126]. A key advantage of VFI lies in its ability to synthesize perceptually smooth and temporally coherent motion, aligning well with the temporal characteristics of the human visual system. HFR content reduces artifacts such as motion blur and judder [86, 42], thereby enhancing the visual quality in high-resolution (HR) and immersive media. This makes VFI particularly valuable in latency-sensitive and fidelity-critical scenarios such as sports broadcasting, interactive gaming, and virtual reality. Finally, in streaming pipelines, VFI also enables bandwidth-efficient video transmission by reconstructing intermediate frames locally, reducing the need to transmit full frame sequences [86].
Figure 1: General process of VFI. Given $M$ consecutive input frames $\{I_{j}\}_{j=0}^{M-1}$, the VFI model $\mathcal{F}$ synthesizes one or more intermediate frames, producing an output sequence $\{\hat{I}_{t_{k}}\}_{k=0}^{N-1}$ with $t_{k}\in(\lfloor\frac{M}{2}\rfloor-1,\lfloor\frac{M}{2}\rfloor)$, where $M\geq 2$ and $N\geq 1$.
As shown in Fig. 1, the general VFI formulation takes a sequence of $M$ consecutive frames $\{I_{j}\}_{j=0}^{M-1}$ to synthesize $N$ intermediate frames $\{\hat{I}_{t_{k}}\}_{k=0}^{N-1}$. The target time indices typically satisfy $t_{k}\in(\lfloor\frac{M}{2}\rfloor-1,\lfloor\frac{M}{2}\rfloor)$. By generating $N$ frames within the central temporal interval, the video frame rate is increased by a factor of $N$. For instance, generating seven intermediate frames per interval transforms a 30fps video into 240fps. We begin by describing the most representative two-input-frame setting ($M=2$). Formally, given two adjacent frames $I_{0}$ and $I_{1}$, a VFI model $\mathcal{F}$ estimates the interpolated frame $\hat{I}_{t}$ at an arbitrary time $t\in(0,1)$:
@@ I Introduction
I-A Methodology Overview
{forest}
Figure 2: Overview of the survey structure.
VFI methodologies can be broadly grouped into classical motion compensation-based [259, 76, 164, 20, 125, 75, 35, 114, 90, 230, 229], deep learning-based [179, 100, 148, 145, 6, 262, 187, 251, 34, 168, 144, 271, 204, 89, 120, 178, 93, 196, 188, 166, 176, 109, 133, 270, 88, 104, 73, 240, 281, 110, 149, 169, 170, 288, 182, 36, 32, 202, 33, 52, 31, 44, 53, 113, 283, 167, 7, 177, 124, 68, 171, 224, 245, 253, 219, 262, 46, 159, 158, 203, 150, 266, 139, 67, 66, 267, 105, 17, 241], and generative frameworks such as diffusion models (DMs) [231, 102, 48, 268, 217, 121, 219, 220, 237, 226, 47, 102, 94, 248, 247, 198, 232, 59, 152, 287, 254, 269, 279, 77, 97, 228, 98, 255, 151, 27, 127, 74, 84]. The motion compensation-based approach dominated the pre-deep-learning era, offering a two-stage strategy: estimating motion explicitly and warping frames accordingly. While effective for simple motion, its reliance on hand-crafted rules and block-based assumptions limits its ability to handle occlusions and complex, non-rigid dynamics. With the advent of convolutional neural networks (CNNs) [122], VFI shifted toward deep learning-based, replacing heuristic pipelines with end-to-end architectures. As a result, they significantly improve robustness under diverse and challenging conditions. Further methodological details are discussed in Sec. II-B. More recently, DMs have been introduced as a generative framework for VFI, framing the task as a conditional denoising process rather than a deterministic frame prediction. This expands the scope of VFI into the emerging concept of Generative Inbetweening [59, 232, 84], enabling uncertainty-aware interpolation and semantically diverse frame synthesis. This shift not only enhances robustness in ambiguous motion scenarios but also opens the door to multi-modal guidance (e.g., text, depth, or motion priors), redefining the role of VFI in creative and interactive video generation.
In parallel to our work, two prior surveys [175, 54] have reviewed VFI techniques, focusing respectively on traditional interpolation methods and early deep learning–based approaches. Compared with these surveys, our paper provides a broader and more up-to-date coverage, including recent Transformer-, Mamba-, and DM-based VFI models that were not available or only briefly discussed in earlier works. Methodologically, we introduce a finer-grained taxonomy and explicitly relates each method to its underlying motion modeling strategy. In addition, we organize existing approaches through a learning-paradigm perspective that distinguishes Center-Time Frame Interpolation (CTFI) from Arbitrary-Time Frame Interpolation (ATFI), and connect these paradigms to their typical training strategies and loss formulations. Finally, our survey systematically compiles VFI datasets and evaluation metrics, providing a practical resource that complements the conceptual taxonomy and goes beyond the scope of previous VFI surveys.
Overview. Fig. 2 shows the overall structure of this paper. Sec. II analyzes methodological taxonomies of VFI. Sec. III introduces and compares the two principal learning paradigms of VFI, and further examines their corresponding training strategies and loss functions. Sec. IV discusses major challenges in VFI, along with how recent methods address them. Sec. V reviews common datasets and evaluation metrics. Section VI explores applications of VFI across diverse domains. Finally, Sec. VII presents future research directions of VFI.
Figure 3: Temporal alignment strategies. Input frames $(I_{0},I_{1})$ or features $(F_{0},F_{1})$ are aligned toward target time $t$ using four strategies: (a) kernel-based, (b) flow-based, (c) attention-based, and (d) cost volume-based. On the right, (a’) (kernel-based), the blue square denotes the fixed kernel support window centered at the output location in $I_{t}$, while the gray patches indicate the actual sampling positions in $I_{t-1}$ and $I_{t+1}$ gathered via learned offsets, so that motion is encoded implicitly through the offset pattern and kernel weights; (b’) (flow-based), the light-blue marks the explicit reference location reached by a displacement vector, showing an explicit motion field; (a’+b’) shows the combined design, where an explicit flow first transports the support window toward a reference region and a local kernel is then applied around that flow-guided position for refinement.
@@ I-A Methodology Overview
Figure 4: General pipeline of VFI. Given two input frames $I_{0}$ and $I_{1}$, deep features $F_{0}$ and $F_{1}$ are first extracted. The features or pixels are then temporally aligned to the target time $t$ using estimated motion, producing $\hat{F}_{0\rightarrow t}$, $\hat{F}_{1\rightarrow t}$ or $\hat{I}_{0\rightarrow t}$, $\hat{I}_{1\rightarrow t}$. A Frame Synthesis module blends the aligned inputs to produce the final frame $\hat{I}_{t}$.
I-B General Pipeline of VFI
@@
The general VFI pipeline consists of four stages: (i) Feature Extraction. Input frames $I_{0}$ and $I_{1}$ are passed through a feature extraction network [207, 191, 39, 244], yielding representations $F_{0}$ and $F_{1}$ for subsequent motion reasoning [173]. (ii) Motion Estimation. Temporal correspondence (i.e., motion) is estimated either explicitly via optical flow [85, 168, 204] or implicitly using learned kernels [169, 170], phase cues [159, 158], attention maps [36, 203], or cost volumes [177, 178]. (iii) Temporal Alignment. The estimated motion is used to temporally align the input pixels or features to the target time $t$, generating $\hat{F}_{0\rightarrow t}$, $\hat{F}_{1\rightarrow t}$ or $\hat{I}_{0\rightarrow t}$, $\hat{I}_{1\rightarrow t}$. As shown in Fig. 3, four major alignment strategies are employed. Although all of them rely on motion estimation, they differ in how motion is represented and applied during alignment. Kernel-based alignment (Fig. 3 (a),(a’)) aggregates local or non-local information from the inputs using learned, spatially-adaptive kernels. These kernels implicitly encode motion by adapting their spatial weights based on local context, allowing motion-aware alignment without explicit flow estimation. As shown in Fig. 3 (a’), the network effectively selects reference locations by assigning larger weights to pixels that contribute most to the output, but the reachable motion range is constrained by the kernel support, making very large displacements harder to capture without excessively increasing the kernel size. Flow-based alignment (Fig. 3 (b),(b’)) warps inputs guided by the estimated flow. Forward warping [168] maps source pixels (i.e., input pixels) to their estimated locations in the target frame. Backward warping [100] samples from the source based on coordinates in the target frame, effectively pulling information from the source toward the desired time. In Fig. 3 (b’), this is visualized as an explicit displacement vector that points from each output query to a reference position in the input frame, which is efficient and well suited to large motion but aggregates information from only a small number of source samples at each step, limiting flexibility under complex or noisy motion. The hybrid kernel-and-flow-based design in Fig. 3 (a’+b’) combines these views by first using an explicit flow vector to transport the sampling window toward an appropriate reference region and then applying a local kernel around that flow-guided position, thereby increasing the number of effective reference samples while preserving the interpretability of the underlying motion field (Sec. II-B3). Attention-based alignment (Fig. 3 (c)) aggregates features based on attention-weighted correspondences [203, 270]. By computing soft correspondences between elements across input frames, this strategy can adaptively focus on semantically relevant regions and align contents even across large spatial-temporal gaps. Cost volume-based alignment (Fig. 3 (d)) constructs dense similarity volumes between feature maps, enabling fine-grained correspondence modeling across space and time. (iv) Frame Synthesis. Finally, the aligned inputs are blended to synthesize the target frame $\hat{I}_{t}$ using simple averaging, weighted blending, or synthesis networks [34] as shown in Fig. 4.
II Methodology
II-A Motion Compensation-based
Before the advent of deep-learning, VFI was mainly tackled via Motion-Compensated Frame Interpolation (MCFI) [164, 75, 35, 90] or Frame Rate Up-Conversion (FRUC) [76, 20, 125, 114, 230, 229]. These approaches, prevalent from the late 1990s through the early 2000s, estimate motion explicitly using block matching or parametric models, then synthesize intermediate frames by warping input frames based on the estimated motion fields.